# One-Shot Imitation under Mismatched Execution

### v1

@@ VI Discussion and Limitations
This work tackles the problem of one-shot imitation from a demonstrator with mismatched execution. We introduce RHyME, a novel framework that leverages task-level correspondences to overcome frame-level visual gaps between the robot and a demonstrator, and learns a video-conditioned policy without access to any paired demonstrations.
Limitations. Although the exact composition of test-time tasks isn’t seen at train-time, transitions between every pair of tasks in the unseen composition must exist in the robot dataset to learn transition actions. This prohibits our method from learning brand-new task sequences in a one-shot fashion.

### v2

@@ VI Discussion and Limitations
This work addresses the challenge of one-shot imitation in the presence of mismatched execution by the demonstrator. We propose RHyME, a novel framework that leverages task-level correspondences to bridge frame-level visual disparities between the robot and the demonstrator, enabling the learning of a video-conditioned policy without paired data.
Limitations. While the exact test-time task compositions are unseen during training, our method relies on transitions between task pairs in the robot dataset to learn transition actions. This limits the ability to learn entirely new task sequences. We note still that our method generalizes well to new compositions when such transitions are present. Expanding the dataset could enhance its generalizability.